OpenStudy (anonymous):

simplify using factoring 96wx^3y^3z^3-144w^3xy^2z^3

OpenStudy (australopithecus):

Example: ab + abc - abc^2 ab(1 + c + c^2) does this help?

OpenStudy (anonymous):

not exactly because im terrible at math

OpenStudy (australopithecus):

Example 2: a^2b^2 + ab^2 + ab^3 ab^2(a + 1 + b)

OpenStudy (australopithecus):

Essentially you are pulling out common multiples from the equation

OpenStudy (australopithecus):

Exponent rule: a^4 = a*a*a*a a^3 = a*a*a a^2 = a*a

OpenStudy (anonymous):

does that include stuff like x^3 and x

OpenStudy (australopithecus):

if you multiply out ab^2(a + 1 + b) = ab^2*a + ab^2*1 + ab^2*b a^2b^2 + ab^2 + ab^3

OpenStudy (australopithecus):

yeah its a general case for any variable or number essentially exponents are an easier way to write out something that is being multiplied out repeatedly

OpenStudy (australopithecus):

so more examples 3^3 = 3*3*3 x^5 = x*x*x*x*x

OpenStudy (anonymous):

so what do i do with the 96w and the 144w^3

OpenStudy (australopithecus):

Negative exponential rule: \[x^{-1} = \frac{1}{x}\] \[\frac{1}{x^{-1}} = x^{1}\] Exponential multiplication rule: \[x^ax^b = x^{a+b}\] \[x^{a}x^{-b} = x^{a + (-b)} = x^{a-b}\] for example: \[x^{3}*x^{5} = x^8\] \[x^2\frac{1}{x^{3}} = x^{2}x^{-3} = x^{2-3} = x^{-1} = \frac{1}{x}\]

OpenStudy (australopithecus):

So for your problem, 96w and 144w^3 by the definition of exponents 144w^3 = 144*w*w*w 144*w*w*w and 96w share a single w, so we can factor it out and get w(144w*w + 96) or w(144w^2 + 96)

OpenStudy (australopithecus):

essentially you are finding common terms and pulling them out

OpenStudy (australopithecus):

Algebra is just a bunch of rules you have to memorize it isn't that tough just look up algebra rules then find practice questions to apply them eventually you wont need the rules you will just get it. Here is a decent list of them http://www.wyzant.com/resources/lessons/math/algebra/properties_of_algebra

OpenStudy (australopithecus):

but yeah back to your question do you want to make an attempt?

OpenStudy (australopithecus):

I can tell you if you are on the right track